OpenStudy (anonymous):

Sue earns $2.00 for each CD she shells and $2.50 for each DVD she sells. Sue earned $950.00 last week selling CD's and DVD s. Write an equation to represent the number of CD's (C) and DVD s(D) Sue sold last week given that she earned $950.00 If Sue sold 305 CD's last week, how many DVD s did she sell?

OpenStudy (anonymous):

she sold 330 DVDs

hero (hero):

Give steps, not answers

OpenStudy (anonymous):

I believe.

OpenStudy (anonymous):

what is the equations

OpenStudy (anonymous):

305c x 2 = 620 $950 - 620 = 330

OpenStudy (anonymous):

Is that good enough for you?

hero (hero):

1. Set up Variables: 305= CDs y = DVDs x = Total Amount of CDs and DVDs 2. Set up Equations: Total Amount Equation: 305 + y = x Total Cost Equation: 2(305) + 2.5y = 950 3. Solve Systems of Equations (a) Original Equations: 305 + y = x 2(305) + 2.5y = 950 (b) Isolate y in both equations: y = x - 305 y = (905 - 610)/2.5 (c) Set y = y and cross multiply: 2.5(x - 305) = 295 (d) Solve for x using order of operations and simplification techniques: 2.5x - 762.50 = 295 2.5x = 1057.5 x = 1057.5/2.5 x = 423 Since x is the total amount of dics sold, then 423 - 305 gives the amount of DVDs sold. 423 - 305 = 118 Therefore 118 DVDs were sold.

OpenStudy (anonymous):

O well your just a smarty pants...

OpenStudy (anonymous):

But she should use mine. lol

hero (hero):

Yours isn't correct unfortunately

OpenStudy (anonymous):

okay. Whatever. I had it checked. By my teacher.

OpenStudy (anonymous):

did u have your checked?

hero (hero):

That just means that you and your teacher are incorrect. You should find a new teacher.

OpenStudy (anonymous):

............. whatever. Think what you want.

OpenStudy (anonymous):

I took Hero's

hero (hero):

lol

OpenStudy (anonymous):

I doubt there was ever a teacher to begin with.

OpenStudy (anonymous):

... Yes. There was.

OpenStudy (anonymous):

@Retract please, keep you Sarcastic comments to your self.

OpenStudy (anonymous):

It wasn't sarcasm.

OpenStudy (anonymous):

@Retract ok what and ever, you can just stop talking to me all around. okay? if you talk to me again I will report you. Have a "good" day